OpenStudy (anonymous):

PLEASE Help!!!!!!!! what is the axis of symmetry for y=4x^2-1 ?

jimthompson5910 (jim_thompson5910):

use the formula x = -b/(2a)

OpenStudy (anonymous):

soo y=-4^2/(1) ?

jimthompson5910 (jim_thompson5910):

4x^2 - 1 is the same as 4x^2 + 0x - 1

OpenStudy (anonymous):

ohh okay so u just solve ?

jimthompson5910 (jim_thompson5910):

in this case, a = 4 and b = 0

jimthompson5910 (jim_thompson5910):

plug those two in and evaluate

OpenStudy (anonymous):

0/4 = 0?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

its not correct .

jimthompson5910 (jim_thompson5910):

well it's x = 0, so there's another typo apparently

jimthompson5910 (jim_thompson5910):

something you'll have to bring up with the teacher

OpenStudy (anonymous):

ohh okay.